### v1

@@ I Introduction
In this work, we consider a single impurity in a strongly correlated 1+1dimensional field theory described by a generalized Gross-Neveu-like $O(3)$ invariant color interaction first studied in [33]. The Gross-Neveu model was studied as a model of elementary particles that is asymptotically free and where dynamical symmetry breaking takes place, giving mass to the fermions [34, 35, 36, 37]. Likewise, the model can also be viewed as a model of superconductivity [38, 39, 40, 41, 42] where an attractive interaction among left and right moving fermions dynamically generates a superconducting mass gap $\Delta=2m$, where $m$ is the mass of a single spinon (kink) excitation.
Notice that in $1+1$ dimensions, spontaneous breaking of a continuous symmetry [43] and the existence of a Goldstone boson is, according to Coleman’s theorem [44], not possible. Thus, the dynamical mass generation in this model is associated with a decoupled O(1) bosonic zero mode, which leads to power decay of the fermionic correlation functions [45].
@@
The model considered in this work can also be viewed as a toy model of QCD Kondo or as a model of the Kondo effect in one-dimensional superconducting wire. Using Bethe Ansatz, we shall provide exact non-perturbative results examining the role of bulk strong interaction on the Kondo effect. We shall return to the $SU(3)$ Gross-Neveu model with impurity as a more realistic description of the QCD Kondo problem in a subsequent paper.
@@
The Kondo effect in superconductors has been immensely studied [25, 24, 26, 46, 47, 48, 49, 50, 51, 52, 53]. A recent work studied a single impurity at the edge of an interacting $SU(2)$ superconducting wire, identifying several impurity phases: the Kondo phase, where the impurity is screened by a multiparticle Kondo cloud; the Yu-Shiba-Rushinov (YSR) phase, where it is screened by a single-particle bound state; and an unscreened phase [2]. Here, we generalize this problem by considering a $SO(3)$ invariant superconducting bulk with spin-$\frac{1}{2}$ impurity in the boundary, and we show that the impurity is overscreened. We shall show, in fact, that for antiferromagnetic coupling, the impurity exhibits four different phases: the Kondo phase where it is overscreened by a many-body Kondo cloud such that the zero temperature entropy of the impurity is $\frac{1}{2}\ln 2$, the zero mode phase where while the impurity is many-body overscreened, there also exists a zero energy boundary excitation that is degenerate with the ground state in the thermodynamic limit, the YSR phase where impurity is unscreened in the ground state but it can be screened by a single particle bound mode formed at the edge of the wire which has finite mass below the bulk mass gap, and finally a local moment phase where the impurity cannot be screened at any energy scale. Apart from the dynamically generated mass gap in the bulk $m$, the model also dynamically generates a Kondo scale $T_{K}$ in the Kondo and zero mode phases that govern the overscreening of the impurity.
@@ II The Model
$H$ $=\int_{-L/2}^{0}\mathrm{d}x\sum_{r,s,u,v=1}^{3}\left[-i(\psi^{\dagger}_{r+}\partial_{x}\psi_{r+}-\psi^{\dagger}_{r-}\partial_{x}\psi_{r-})+2g\,\psi^{\dagger}_{r+}\psi^{\dagger}_{s-}(\sum_{\alpha}\tau^{\alpha}\otimes\tau^{\alpha}+I)_{rs}^{uv}\psi_{u-}\psi_{v+}+2J\,\psi^{\dagger}_{r,-}(0)\vec{\tau}_{rs}\psi_{s,+}(0)\cdot\vec{\sigma}\right]$ (1)
@@
$I_{ac}^{bd}=\delta_{a,b}\delta_{c,d}$ is the identity matrix, $\vec{\sigma}$ are the $SU(2)$ Pauli matrices acting on the spin space of the spin-$\frac{1}{2}$ impurity localized at $x=0$, $\pm$ denotes the chirality and $r,s=1,2,3$ is the $O(3)$ color index. The coupling $g>0$ describes the bulk interactions of both the attractive spin and charge degrees of freedom between the fermions of opposite chirality, and the coupling $J>0$ describes the chirality-changing interaction between the bulk fermion and the localized impurity at the right edge of the chain.
When $J=0$, the model reduces to Bethe Ansatz solvable $O(3)$ invariant model considered in [33] where due to the dynamical symmetry breaking the model opens a mass gap and contains fractionalized kink excitations. Likewise, when $g=0$, the model reduces to Bethe Ansatz solvable case of the higher spin Kondo model studied in [54] where the spin-$\frac{1}{2}$ impurity is overscreened by spin-$1$ bulk and the exhibits behavior similar to the two-channel Kondo physics [55]. Here, we shall show that when both $J$ and $g$ are non-zero, the model still exhibits exact solution via Bethe Ansatz [56, 57, 58, 59, 60, 61, 62, 63]. Note that due to isomorphism $\mathfrak{su}(2)\cong\mathfrak{so}(3)$, it is also possible to view the bulk Hamiltonian as $SU(2)$ spin-$1$ Gross-Neveu model (see Appendix A). Since the boundary interaction is with spin-$\frac{1}{2}$ spin, it is often useful to use $SU(2)$ representation to label the states in the theory.
Figure 1: The weak-coupling RG flow diagram for Hamiltonian Eq.(1) are given by Eq.(4) and Eq.(5). The flow is away from the non-interacting point $J=0=g$. The orange curves denote the overscreened Kondo phase, where the boundary coupling flows to a non-trivial 2-channel Kondo fixed point. The purple lines represent the zero-mode phase, where the boundary coupling flows to the 2-channel fixed point, but there is a boundary excitation of zero energy. Magenta lines indicate the YSR phase, where the impurity is unscreened in the ground state but screened by single-particle bound modes in excited states. Cyan lines depict the unscreened regime, where the boundary coupling flows to zero such that impurity is unscreened. The red curves demarcate the four boundary phases of the model.
When the boundary coupling $J$ is zero, the bulk model was proven to be integrable in [33] and studied in depth using both large $N$ analysis and integrability technique. Here, we shall explicitly construct the eigenstate of the Hamiltonian Eq.(1) in Appendix B and then prove that the model is still integrable in the presence of the boundary perturbation in Appendix C. We shall derive the Bethe Ansatz equations of the model using the functional Bethe Ansatz method [64, 65] and fusion hierarchy [65, 66, 67, 68] in Appendix C and alternatively by using the fusion technique of [33] in Appendix D.
@@
obtained from the exact solution in Appendix F in an unconventional cutoff scheme where the UV cut-off is imposed in the fully interacting theory [14]. Since $\beta(g)$ is always negative, $g$ always flows to the strong coupling where the Hamiltonian Eq.(1) gains mass via dimensional transmutation. In contrast, $\beta(J)$ can change sign depending on the relative strengths of $J$ and $g$; therefore $J$ flows to strong coupling where the impurity is overscreened by many body Kondo cloud only when $J\gg g$, but for $J\ll g$ it flows back to weak coupling, as shown in Fig. 1, and the impurity is not screened. In intermediate regimes where $g\sim J$, there exist two phases. One we call the zero mode phase, where the impurity is still overscreened by the many-body cloud, but a unique zero-energy boundary excitation is given by an isolated boundary string solution of the Bethe Ansatz equations. Further, there exists another phase that we call the YSR phase  [^1] Similar to the Yu-Shiba-Rushinov phase in BCS superconductors, where the impurity is no longer screened in the ground state, but there is elementary boundary excitation where a single particle in an edge localized bound mode screens the impurity. The phases correspond to definite ranges of the RG invariant $d(J,g)$
@@
that remains constant while both $g$ and $J$ are running coupling constants that vary with the energy scale. This parameter could take either real value or purely imaginary value. In Fig.1, constant $d$ surfaces are shown via lines with arrows and denote the boundaries between the phases. These phases are plotted in Fig.1. We shall also use this parameter $d$ to label various impurity phases in the model in the later sections. Notice that, at the diagonal line $J=2g$, the two equations (Eq.(4) and Eq.(5)) become the same, where the flow lines above represent the Kondo phase, whereas the lines below represent the other three phases.
@@ III Summary of results
Figure 2: The phase diagram of the spin-$1$ Gross-Neveu model with a Kondo impurity with spin $S=\frac{1}{2}$ at the boundary illustrates the spectrum’s dependence on $d$, the RG invariant parameter that combines both the bulk and boundary running coupling constants. In the Kondo regime, $d\in\mathbb{R}$ or $20<\delta<1/$ 2 ($\delta=id$), where the Kondo scale $T_{K}$ greater than $m$ (The dynamically generated mass of the spinon) the impurity is overscreened by a multiparticle Kondo cloud. For the parametric range $11/2<\delta<$ 1, while the multiparticle Kondo cloud still overscreens the impurity, there exists a unique zero-energy excitation (shown by the solid magenta lines) in the thermodynamic limit which is described by a unique purely imaginary solution of the Bethe Ansatz equation $\lambda=\pm i\left(\frac{1}{2}-\delta\right)$. In the range, $1<\delta<{2}$, the YSR phase, the purely imaginary solution, acquires finite mass $m\cos(\pi\delta)$ i.e. corresponding to mid-gap states with energy shown in the solid magenta curve. In this phase, the impurity is no longer screened by the multiparticle cloud in the ground state but by a single particle bound state exponentially localized at the edge of the system. Finally, for $\delta>{2}$, the impurity remains unscreened, and the mid-gap states are absent. The boundary string solution has zero energy in this regime, and it is not possible to add this solution to the ground state without adding an additional even number of holes [2]. The orange-shaded region denotes the continuum of excitation above the mass gap $\Delta=2m$. Notice that this result is in sharp contrast with the classical impurity in the BCS superconductor, where the YSR phase exists in the entire phase space. The phase diagram shares similarities with the spin-1/2 Gross-Neveu model with a single Kondo impurity studied in [27, 2] where the Kondo, YSR, and the unscreened phases appear, and the bound state in the YSR regime could have negative energy in the ground state.
The bulk term of the model Hamiltonian Eq.(1) describes a one-dimensional superconductor where the superconducting mass gap $\Delta=2m$ is dynamically generated and the two-point function of the fermionic bilinear $\psi^{\dagger}_{a-}\psi_{a+}(x)$ acquires non-vanishing vacuum expectation value that falls off algebraically, thereby showing quasi-long range order [33]. When a spin-$\frac{1}{2}$ particle is attached to the edge of the superconductor, it is the interplay between the bulk attractive interaction and the boundary Kondo effect that leads to an interesting impurity behavior. In terms of an RG invariant parameter $d$ that combines both the bulk and boundary running coupling constants, the boundary physics can be classified into four distinct phases as shown in the boundary phase diagram Fig.2. When the parameter $d\in\mathbb{R}$ or when $d=i\delta$ and $0<\delta<\frac{1}{2}$, corresponding to large impurity coupling $J$ compared to the bulk interaction strength $g$ the impurity is in an overscreened Kondo phase where a multiparticle cloud of massive spinons in the bulk overscreens the spin-1/2 impurity. A new Kondo scale $T_{K}$ is dynamically generated in this phase, which, among others, characterizes the overscreening cloud. When $\frac{1}{2}<\delta<1$, an intermediate phase appears where the impurity is still dynamically overscreened by the multiparticle cloud, but there is a unique boundary excitation where a boundary localized edge mode and a one-string solution of Bethe equations form a zero energy singlet excitation. So far, when the RG invariant parameter $d$ is real or when $d=i\delta$ is between $10<\delta<$ 1, the impurity is screened in the ground state by a multiparticle Kondo cloud. At $\delta=1$, the model undergoes boundary quantum phase transition where for $\delta>1$, the impurity is no longer screened in the ground state. However, when $21<\delta<$ 2, a mid-gap state forms that can screen the impurity, albeit not in the ground state  [^2] We note that an analogous mid-gap bound mode would exist throughout the phase space in a BCS superconductor with classical impurity [25, 24, 26] and screen it in the ground state. Here, however, the strong quantum fluctuation restricts it to a small region in the phase space. Finally, when $\delta>2$, the bulk interaction strength is larger than the boundary interaction strength, and neither the Kondo physics is stabilized nor the midgap state survives in the spectrum. Thus, the impurity is no longer screened on any scale, and the model is in the local moment phase.

### v2

@@ I Introduction
We begin by studying a single impurity in a strongly correlated 1+1-dimensional field theory described by an attractive 4-fermion $O(3)$ invariant color interaction first studied in [33]. The O(3) model generalizes the standard SU(2) Gross-Neveu model that was introduced as a model of elementary particles that is asymptotically free and where dynamical symmetry breaking takes place, giving mass to the fermions [34, 35, 36, 37]. Likewise, the model can be viewed as a model of superconductivity [38, 39, 40, 41, 42] where the attractive interaction among left and right-moving fermions dynamically generates a superconducting mass gap $\Delta=2m$, where $m$ is the mass of a single spinon (kink) excitation. However, since in $1+1$ dimensions the spontaneous breaking of a continuous symmetry [43] and the existence of a Goldstone boson is not possible, according to Coleman’s theorem [44], the dynamical mass generation in this model is associated with a decoupled O(1) bosonic zero mode, which leads to power decay of the fermionic correlation functions [45].
@@
The model considered in this work can also be viewed as a toy model of QCD Kondo or as a model of the Kondo effect in a one-dimensional superconducting wire. Using Bethe Ansatz, we shall provide exact non-perturbative results examining the role of bulk strong interaction on the Kondo effect. We address the $SU(3)$ Gross–Neveu model with impurity as a realistic description of the QCD Kondo problem in Ref. [46].
The Kondo effect in superconductors has been widely studied [25, 24, 26, 47, 48, 49, 50, 51, 52, 53, 54]. In a superconductor, electrons interact attractively to form Cooper pairs throughout the bulk, establishing a coherent superconducting state. However, when an impurity is present, the conduction electrons also interact with the impurity to form a many-body Kondo singlet, and further, when the interaction with the impurity is strong enough, a local bound state may be formed that screens the impurity. The competition between the formation of Cooper pairs and the screening of the impurity by a singlet state leads to unexpected phenomena, including unconventional superconductivity and quantum critical behavior. A recent work studied a single impurity at the edge of an interacting $SU(2)$ superconducting wire and identified several impurity phases: the Kondo phase, where the impurity is screened by a multiparticle Kondo cloud; the Yu-Shiba-Rusinov (YSR) phase, where it is screened by a single-particle bound state; and an unscreened phase [2].
@@
Here, we generalize this problem by considering a $SO(3)$ invariant superconducting bulk interacting with spin-$\frac{1}{2}$ impurity in the boundary, and we show that the impurity is overscreened. We shall show, further, that for antiferromagnetic coupling, the impurity exhibits four different phases: i. the Kondo phase where it is overscreened by a many-body Kondo cloud where the zero temperature entropy of the impurity is $\frac{1}{2}\ln 2$, ii. the zero mode phase where the impurity is many-body overscreened and there also exists a zero energy boundary excitation so that the ground state is degenerate with the ground state in the Kondo phase in the thermodynamic limit, iii. the YSR phase, where the impurity is unscreened in the ground state but can be screened by a single particle bound mode formed at the edge of the wire, which has a finite mass below the bulk mass gap, and finally iv. a local moment phase, where the impurity cannot be screened at any energy scale. Apart from the dynamically generated mass gap in the bulk $m$, the model also dynamically generates a Kondo scale $T_{K}$ in the Kondo and zero mode phases that govern the overscreening of the impurity.
@@ II The Model
$H$ $=\int_{-L}^{0}\mathrm{d}x\sum_{r,s,u,v=1}^{3}\left[-i(\psi^{\dagger}_{r+}\partial_{x}\psi_{r+}-\psi^{\dagger}_{r-}\partial_{x}\psi_{r-})+2g\,\psi^{\dagger}_{r+}\psi^{\dagger}_{s-}(\sum_{\alpha}\tau^{\alpha}\otimes\tau^{\alpha}+I)_{rs}^{uv}\psi_{u-}\psi_{v+}+2J\,\psi^{\dagger}_{r,-}(0)\vec{\tau}_{rs}\psi_{s,+}(0)\cdot\vec{\sigma}\right],$ (1)
@@
$I_{ac}^{bd}=\delta_{a,b}\delta_{c,d}$ is the identity matrix, $\vec{\sigma}$ are the $SU(2)$ Pauli matrices acting on the spin space of the spin-$\frac{1}{2}$ impurity localized at $x=0$, $\pm$ denotes the chirality and $r,s=1,2,3$ is the $O(3)$ color index. The coupling $g>0$ describes the bulk interactions of both the attractive spin and charge degrees of freedom between the fermions of opposite chirality and the coupling $J>0$ describes the chirality-changing interaction between the bulk fermion and the localized impurity at the right edge of the chain.
When the boundary coupling $J$ is zero, the bulk model reduces to a Bethe Ansatz-solvable $O(3)$ [33] and was studied in depth using both large $N$ analysis and integrability techniques. It was shown that a dynamical symmetry breaking takes place, and the model opens a mass gap and contains fractionalized kink excitations. On the other hand, when $g=0$, the model reduces to the Bethe Ansatz solvable case of the higher spin Kondo model studied in [55] where the spin-$\frac{1}{2}$ impurity is overscreened by spin-$1$ bulk and exhibits behavior similar to the two-channel Kondo found in [56].
Here, we shall show that when both $J$ and $g$ are non-zero, the model can still be solved exactly via the Bethe Ansatz approach  [57, 58, 59, 60, 61, 62, 63, 64]. We shall explicitly construct the eigenstate of the Hamiltonian Eq.(1) in Appendix B and then prove that the model is still integrable in the presence of the boundary perturbation in Appendix C. We shall derive the Bethe Ansatz equations of the model using the functional Bethe Ansatz method [65, 66] and fusion hierarchy [66, 67, 68, 69] in Appendix C and alternatively by using the fusion technique of [33] in Appendix D.
Note that due to isomorphism $\mathfrak{su}(2)\cong\mathfrak{so}(3)$, it is also possible to view the bulk Hamiltonian as $SU(2)$ spin-$1$ Gross-Neveu model (see Appendix A). Since the boundary interaction is with spin-$\frac{1}{2}$ spin, it is often useful to use $SU(2)$ representation to label the states in the theory.
@@
obtained from weak-coupling perturbation theory or from the exact solution in Appendix F  [^1] The exact RG equations obtained from the Bethe Ansatz solution are expressed in an unconventional cutoff scheme where the UV cut-off is imposed in the fully interacting theory. However, at weak coupling, both schemes coincide [14].. Since $\beta(g)$ is always negative, $g$ always flows to the strong coupling where the Hamiltonian Eq.(1) gains mass via dimensional transmutation. In contrast, $\beta(J)$ can change sign depending on the relative strengths of $J$ and $g$; therefore $J$ flows to strong coupling where the impurity is overscreened by many body Kondo cloud only when $J\gg g$, but for $J\ll g$ it flows back to weak coupling, as shown in Fig. 1, and the impurity is not screened. In intermediate regimes where $g\sim J$, there exist two phases. One we call the zero mode phase, where the impurity is still overscreened by the many-body cloud, but a unique zero-energy boundary excitation is given by an isolated boundary string solution of the Bethe Ansatz equations. Further, there exists another phase that we call the YSR phase, which is similar to the Yu-Shiba-Rusinov phase in BCS superconductors, where the impurity is no longer screened in the ground state, but there is an elementary boundary excitation where a single particle in an edge-localized bound mode screens the impurity.
The phases correspond to definite ranges of the RG invariant $d(J,g)$
@@
that remains constant while both $g$ and $J$ are running coupling constants that vary with the energy scale. This parameter could take either a real value or a purely imaginary value. In Fig.1, constant $d$ surfaces are shown via lines with arrows and denote the boundaries between the phases. These phases are illustrated in Fig.1. We shall also use this parameter $d$ to label various impurity phases in the model in the later sections. Notice that, at the diagonal line $J=2g$, the two equations (Eq.(4) and Eq.(5)) become the same, where the flow lines above represent the Kondo phase, whereas the lines below represent the other three phases. The weak-coupling renormalization group equations (Eqs. (4) and (5)) (obtained exactly from the Bethe Ansatz analysis) describe how the couplings evolve from the non-interacting point ($g=J=0$). They reveal that, depending on the relative strength of $g$ and $J$, the boundary coupling $J$ can either grow strong or diminish to zero, indicating at least two different regimes. By diagonalizing the Hamiltonian and carefully analyzing the spectrum, we find that the full-phase diagram actually comprises four distinct boundary phases, each exhibiting a unique low-energy behavior. Although the RG equations reveal how the two couplings evolve, they do not determine whether the model flows to strong or weak coupling fixed points (if such fixed points exist). To pinpoint the infrared fixed points, we employ the Bethe Ansatz.
Figure 1: The weak-coupling RG flow diagram for Hamiltonian Eq.(1) is given by Eq.(4) and Eq.(5). The flow is away from the non-interacting point $J=0=g$. The blue curves denote the overscreened Kondo phase, where the boundary coupling flows to a non-trivial 2-channel Kondo fixed point. The purple lines represent the zero-mode phase, where the boundary coupling flows to the 2-channel fixed point, but there is a boundary excitation of zero energy. Pink lines indicate the YSR phase, where the impurity is unscreened in the ground state but screened by single-particle bound modes in excited states. Green lines depict the unscreened regime, where the boundary coupling flows to zero, such that the impurity is unscreened. The dashed black curves demarcate the four boundary phases of the model.
@@ III Summary of results
The bulk term of the model Hamiltonian Eq.(1) possesses several interesting properties such as dimensional transmutation, asymptotic freedom, and fractionalized excitations  [37]. It describes a one-dimensional superconductor where the superconducting mass gap $\Delta=2m$ is dynamically generated and the two-point function of the fermionic bilinear $\psi^{\dagger}_{a-}\psi_{a+}(x)$ acquires a non-vanishing vacuum expectation value that falls off algebraically, thereby showing quasi-long-range order [33]. When a spin-$\frac{1}{2}$ impurity is attached to the edge of the superconductor, it is the interplay between the bulk attractive interaction (that leads to superconducting order) and the boundary Kondo effect (where cloud of bulk fermion overscreens the impurity) that leads to a very rich impurity behavior: Both the superconducting coupling $g$ and the Kondo coupling $J$ are running coupling constants. In the absence of the Kondo coupling $J$, the superconducting coupling $g$ flows to the strong coupling and opens a mass gap in the spin sector of the theory, while the charge decouples and remains massless. Likewise, in the absence of the superconducting coupling $J$, the Kondo coupling $g$ flows to an intermediate non-Fermi liquid fixed point where the impurity is overscreened by the bulk fermions such that there is residual impurity entropy of $\frac{1}{2}\ln 2$ signaling a fractionalized degree of freedom (Majorana) at zero temperature. When both bulk superconducting coupling $g$ and the boundary Kondo coupling $J$ are present, they both flow. In such a case, we identify an RG invariant parameter $d$ that combines both the bulk and boundary running coupling constants. In terms of the RG invariant parameter $d$, the boundary physics can be classified into four distinct phases as shown in the boundary phase diagram Fig. 2. When the parameter $d\in\mathbb{R}$ or when $d=i\delta$ and $0<\delta<\frac{1}{2}$, corresponding to large impurity coupling $J$ compared to the bulk interaction strength $g$ the impurity is in an overscreened Kondo phase where a multiparticle cloud of massive spinons in bulk overscreens the spin-1/2 impurity. A new Kondo scale $T_{K}$ is dynamically generated in this phase, which, among others, characterizes the overscreening cloud. Moreover, there is residual impurity entropy of $\frac{1}{2}\ln 2$ just like in the conventional two-channel Kondo problem. When $\frac{1}{2}<\delta<1$, an intermediate phase appears where the impurity is still dynamically overscreened by the multiparticle cloud, but there is a unique boundary excitation where a boundary localized edge mode and a one-string solution of Bethe equations form a zero energy singlet excitation. This phase blends features of the Kondo phase (impurity screening) and the YSR phase (presence of a boundary excitation). So far, when the RG invariant parameter $d$ is real or when $d=i\delta$ is between $10<\delta<$ 1, the impurity is screened in the ground state by a multiparticle Kondo cloud. At $\delta=1$, the model undergoes a boundary quantum phase transition, where for $\delta>1$, the impurity is no longer screened in the ground state. However, when $21<\delta<$ 2, a mid-gap state forms that can screen the impurity, albeit not in the ground state. Notice that an analogous mid-gap bound mode would exist throughout the phase space in a BCS superconductor with classical impurity [25, 24, 26] and screen it in the ground state. Here, however, the strong quantum fluctuation restricts it to a small region in the phase space. Finally, when $\delta>2$, the bulk interaction strength is larger than the boundary interaction strength, and neither the Kondo physics is stabilized nor the midgap state survives in the spectrum. Thus, the impurity is no longer screened on any scale, and the model is in the local moment phase.
Phases Parameter range Properties
Overscreened Kondo $d\in\mathbb{R}\cup\delta\in(0,1/2)$ Impurity is overscreened by the multiparticle cloud.
No boundary excitations.
Kondo screening wins over the superconducting order.
Zero mode $\delta\in(1/2,1)$ Impurity is overscreened by the multiparticle cloud.
There exists a zero-energy boundary excitation.
YSR $\delta\in(1,2)$ Impurity is not screened in the ground state.
Impurity is screened in the mid-gap state.
Finite energy boundary excitations.
Local moment $\delta>2$ Impurity is not screened at any energy scale.
Superconducting order wins over the Kondo screening.
Table 1: Four impurity phases and their key distinguishing properties.
In summary, when the bulk interaction $g$ is absent, the model exhibits two-channel Kondo behavior throughout the phase for antiferromagnetic Kondo coupling $J$, where the impurity is overscreened by the Kondo screening cloud. Likewise, when the Kondo spin exchange interaction $J$ is absent, the model exhibits superconducting behavior where the attractive spin density-spin density interaction in the bulk opens a spin gap. When both the couplings are present the relative strengths of the two couplings lead to a rich boundary phenomenon. When $J\gg g$, the bulk exhibits superconducting order while at the boundary, the impurity is screened by a multiparticle Kondo cloud, just like in the conventional two-channel Kondo model. On the opposite limit, when $g\gg J$, the bulk superconductivity persists, but at the boundary, the Kondo impurity is entirely unscreened as if the Kondo coupling were ferromagnetic, which can be understood from the fact that the sign of the beta function in the flow equation Eq.(5) changes to positive. There exist two intermediate phases when $g\sim J$ where in both cases the bulk has superconducting order, but in one case, the impurity is overscreened by a multiparticle cloud and in addition there exists a zero-energy boundary excitation,n whereas in another case, the impurity is unscreened in the ground state but there exists a mid-gap state where the impurity is screened by this single-particle state. We briefly tabulate the main distinguishing features between the four phases in Table 1
@@
Figure 2: The phase diagram of the spin-$1$ superconductor with a spin $S=\frac{1}{2}$ Kondo impurity at the boundary. It shows the spectrum dependence on $d$, the RG invariant parameter that combines both the bulk and boundary running coupling constants. In the Kondo regime, $d\in\mathbb{R}$ or $20<\delta<1/$ 2 ($\delta=id$), where the Kondo scale $T_{K}$ is greater than $m$ (The dynamically generated mass of the spinon), the impurity is overscreened by a multiparticle Kondo cloud. For the parametric range $11/2<\delta<$ 1, while the multiparticle Kondo cloud still overscreens the impurity, there exists a unique zero-energy excitation (shown by the solid magenta lines) in the thermodynamic limit which is described by a unique purely imaginary solution of the Bethe Ansatz equation $\lambda=\pm i\left(\frac{1}{2}-\delta\right)$. In the range, $1<\delta<{2}$, the YSR phase, the purely imaginary solution, acquires finite mass $m\cos(\pi\delta)$ i.e. corresponding to mid-gap states with energy shown in the solid magenta curve. In this phase, the impurity is no longer screened by the multiparticle cloud in the ground state but by a single particle bound state exponentially localized at the edge of the system. Finally, for $\delta>{2}$, the impurity remains unscreened, and the mid-gap states are absent. The boundary string solution has zero energy in this regime, and it is not possible to add this solution to the ground state without adding an additional even number of holes [2]. The orange-shaded region denotes the continuum of excitation above the mass gap $\Delta=2m$. Notice that this result is in sharp contrast with the classical impurity in the BCS superconductor, where the YSR phase exists in the entire phase space. The phase diagram shares similarities with the spin-1/2 Gross-Neveu model with a single Kondo impurity studied in [27, 2] where the Kondo, YSR, and the unscreened phases appear, and the bound state in the YSR regime could have negative energy in the ground state.